brettandadrionna:

The function y=h(x) is defined as follows:

brettandadrionna:

brettandadrionna:

@Five @MrMudd183

Five:

hey

Five:

Choices?

brettandadrionna:

the answer is h(1)_____ and h(4)______ @Five

brettandadrionna:

Five:

ooo @jhonyy9 @justus help me out ??

senpai:

@Five ima big fan

ajh:

@Five ima big fan

ExoticWinter:

@Five ima big fan

SnowSzn:

@Five ima big fan

hiiiiiiiiiiiiii:

@Five ima big fan

JUNIOR4256:

@Five ima big fan

Five:

Fans plz stop

justus:

one moment

justus:

When x is 1 and the function is h(x)=x^22 then it will be h(1)=(1)^2=1 When x=4 the function is h(x)=5 So h(4)=5 h(1)=1 h(4)=5

justus:

x^2 *

Five:

thanks Justus

justus:

Of course.

justus:

so the answers would be h(1)=1, h(4)=5

loyaltiesleah:

i love you @Five

youandme:

@Five ima big fan

brettandadrionna:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @loyaltiesleah i love you @Five \(\color{#0cbb34}{\text{End of Quote}}\) unrelated

JUNIOR4256:

no there fans

brettandadrionna:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @justus so the answers would be h(1)=1, h(4)=5 \(\color{#0cbb34}{\text{End of Quote}}\) thanks

justus:

yep